10) Please calculate the Sensitivity and Specificity from the 2 by 2 (Screening Test A). (2 points) Test + Test- Diseased 100 10 Not Diseased 20 80 Sensitivity Specificity How many false positives are there? (1 point)

12) Please Calculate the Sensitivity and Specificity of Screening Test B (2 points) Not D Test Positive 34 Test Negative 13 If I wanted to limit the number of false positives, which of the two screening tests would be preferred? A or B? (1 point) 13) 20 People have disease and 80 people do not have a disease. The sensitivity of the test is 0.75. The specificity of the test is 0.60. Fill in the following table ( points) Disease Present No Disease Present Total Test+ Test -
Answer #1

Solution :- Given that,

Sensitivity = (100/(100+10)) = 100/110= 0.909

Specificity = (80/(80+20)) = 80/100 = 0.80